OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with a first term of 2 and a second term of -8?

OpenStudy (skyz):

Beep Boop Beep Beep bOOP

OpenStudy (anonymous):

>o>

OpenStudy (skyz):

The calculator ndoesnt like ur problem...

OpenStudy (anonymous):

lmfao calculator nvr helps anymoar

OpenStudy (skyz):

>.<

OpenStudy (anonymous):

The explicit equation is \[a _{n}=a _{1}* q ^{n-1}\]In your case \[a _{n}= 2 * (-4)^{n-1}\] or to put it more nicely: \[a _{n}= 2 * \frac{ (-4)^{n} }{ 4 } = \frac{ (-4)^{n} }{ 2 }\] (q is -4 because you divide the second term by the first) I'm not sure about the domain though. I tried finding the minimum/maximum but the expression I got was \[\frac{ (-4)^{n} }{ 2 } + \frac{ (-4) ^{n}}{ 8 }\] which is positive for all even n-s and negative for odd ones.

OpenStudy (anonymous):

a. an = 2(-8)n - 1; all integers where n ≥ 1 b. an = 2(-8)n - 1; all integers where n ≥ 0 c. an = 2(-4)n - 1; all integers where n ≥ 0 d. an = 2(-4)n - 1; all integers where n ≥ 1 uhhh these r the options >_< im thinking mayb a?

OpenStudy (anonymous):

It's d because it would be a rational number if you plugged in 0 for n, so it has to be greater than or equal to 1.

OpenStudy (anonymous):

okay :D thank you

OpenStudy (anonymous):

glad to help :)